OpenStudy (anonymous):

what percent of 96 is 72? WILL GIVE MEDAL AND FANS :)

OpenStudy (anonymous):

I have a couple more after this one so if you want more medals I will post them seperately :)

OpenStudy (lizz123):

Set the equation up like this, "is" always represents the equals sign, "what" is the variable: 72 = x * 96 Divide both sides of the equation by 96 to make the variable stand alone: 72/96 = 96x/96 Leaving you with: .75 = x Now, multiply .75 by 100 to get the answer in percentage form: 75%

OpenStudy (anonymous):

okay I am going to go post the next question.

OpenStudy (lizz123):

k